sqrt 2x+7 - sqrt x+7 = 1

Answer by stanbon(75887)

sqrt 2x+7 - sqrt x+7 = 1
----
sqrt(2x+7) = 1+sqrt(x+7)
Square both sides to get:
2x+7 = 1 + 2sqrt(x+7) + x+7
-----
x - 1 = 2sqrt(x+7)
Square both sides
----
x^2 -2x + 1 = 4x + 28
-----
x^2 - 6x -27 = 0
----
x = -3 or x = 9
-----
Check each of those answers in the original equation.
-------------------
Cheers,
Stan H.